# Microscopic Variability Alters Macroscopic Rotation Speed in Stochastic Spiral Waves

### v1

@@ I Introduction
In this work, we investigate the impact of noise on spiral waves. By introducing controlled perturbations, we analyze how noise modifies the core properties of these waves, leading to shifts in their trajectories and changes in their rotation speed. We employ complementary analytical and numerical approaches: the analytical framework decomposes the noise-induced correction into instantaneous and orbital-drift contributions, while in silico simulations of the Barkley model provide direct verification and mutual validation. Together, these methods yield a coherent picture of how stochasticity reshapes spiral-wave dynamics in excitable media.
@@ II.2 Validation
Additive and attenuating temporal noise give values of 11.77 and 5.89 for $\omega^{(2)}$, and -0.69 and -0.61 for $\omega^{(2)}_{\rm od}$ respectively. Hence, while the instantaneous term $\omega^{(2)}$ always causes rotational speed slowdown, the orbital drift term $\omega^{(2)}_{\rm od}$ can both accelerate or slow down rotation.
@@ III Discussion
Our results further show that the influence of noise depends on its type: spatiotemporal noise is the most robust, allowing large amplitudes before propagation failure or spiral breakup occurs, while purely spatial or temporal noise destabilizes waves at much lower noise levels (Fig. 3). However, noise has a larger influence on the slowdown of purely spatial or temporal noise than it has on spatiotemporal noise. This distinction has direct implications for cardiac tissue, where natural spatiotemporal fluctuations may shift toward predominantly spatial noise with aging or fibrosis [dhein2021remodeling, panfilov2002spiral]. In such cases, spirals may initially slow further, but when the noise level rises too high, spiral breakup becomes inevitable. Importantly, static spatial heterogeneity is already known to cause conduction slowing [nezlobinsky2020anisotropic, de2018arrhythmogenicity] and break-up, albeit before a spiral has formed. However, once a spiral has formed, our findings suggest that in addition to this baseline effect, stochastic perturbations can further stabilize and slow spiral waves.
@@ Appendix B Noise realization
For temporal noise, the function $\xi$ in Eq. (18) is sampled from a zero-mean unit-variance normal distribution at each timestep. In this case $Q$ and $\mathrm{Tr}$ both reduce to identity operators. For spatial-temporal noise, one chooses an orthonormal basis $(e_{k})_{k\geq 0}$ for the space of square-integrable functions, picks a spatial covariance function $q$ and writes Qv=qvQv=q*v for the associated convolution operator. For white noise one substitutes the delta function for $q$. In Eq. (18) one uses the sum $\xi=\sqrt{Q}e_{k}\xi_{k}$ where each $\xi_{k}$ is again sampled from a zero-mean unit-variance normal distribution at each time. Spatial noise is sampled in the same fashion at $t=0$ but remains constant thereafter, while $\sqrt{dt}$ is replaced by dtdt in Eq. (18).

### v2

@@ I Introduction
In this work, we investigate the impact of noise on spiral waves. By introducing controlled perturbations, we analyze how noise modifies the core properties of these waves, leading to shifts in their trajectories and changes in their rotation speed. We employ complementary analytical and numerical approaches: the analytical framework decomposes the noise-induced correction into instantaneous and orbital-drift contributions, while in silico simulations of the Barkley model provide direct verification and mutual validation. Together with recent stability results [kuehn2022stochastic], these methods yield a coherent picture of how stochasticity reshapes spiral-wave dynamics in excitable media.
@@ II.2 Validation
Additive and attenuating temporal noise give values of -11.77 and -5.89 for $\omega^{(2)}$, and 0.69 and 0.61 for $\omega^{(2)}_{\rm od}$ respectively. Hence, while the instantaneous term $\omega^{(2)}$ always causes rotational speed slowdown, the orbital drift term $\omega^{(2)}_{\rm od}$ can both accelerate or slow down rotation.
@@ III Discussion
Our results further show that the influence of noise depends on its type: spatiotemporal noise is the most robust, allowing large amplitudes before propagation failure or spiral breakup occurs, while purely spatial or temporal noise destabilizes waves at much lower noise levels (Fig. 3). However, noise has a larger influence on the slowdown of purely spatial or temporal noise than it has on spatiotemporal noise. This distinction has direct implications for cardiac tissue, where natural spatiotemporal fluctuations may shift toward predominantly spatial noise with aging or fibrosis [dhein2021remodeling, panfilov2002spiral]. In such cases, spirals may initially slow further, but when the noise level rises too high, spiral breakup becomes inevitable. Importantly, static spatial heterogeneity is already known to cause conduction slowing [nezlobinsky2020anisotropic, de2018arrhythmogenicity] and break-up, albeit before a spiral has formed. Once a spiral has formed, our findings suggest that in addition to this baseline effect, stochastic perturbations can further stabilize and slow spiral waves.
@@ Acknowledgements.
Data Availability
The data that support the findings of this article are openly available [CodeStochasticSpirals].
@@ Appendix B Noise realization
For temporal noise, the function $\xi$ in Eq. (18) is sampled from a zero-mean unit-variance normal distribution at each timestep. In this case $Q$ and $\mathrm{Tr}$ both reduce to identity operators. For spatial-temporal noise, one chooses an orthonormal basis $(e_{k})_{k\geq 0}$ for the space of square-integrable functions, picks a spatial covariance function $q$ and writes Qv=qvQv=q*v for the associated convolution operator. For white noise one substitutes the delta function for $q$. In Eq. (18) one uses the sum $\xi=\sqrt{Q}e_{k}\xi_{k}$ where each $\xi_{k}$ is again sampled from a zero-mean unit-variance normal distribution at each time. Spatial noise is sampled in the same fashion at $t=0$ but remains constant thereafter, while $\sqrt{dt}$ is replaced by $dt$ in Eq. (18).